# The Impact of Partial Computations on the Red-Blue Pebble Game

### v1

@@ 6.3.2. Matrix Multiplication
Matrix multiplication is a fundamental operation in computer science, and also an essential building block in various applications. Given two matrices of sizes $m_{1}\times m_{2}$ and $m_{2}\times m_{3}$, respectively, we consider the standard matrix multiplication algorithm which executes $m_{1}\!\cdot\!m_{2}\!\cdot\!m_{3}$ multiplications. The I/O-efficient execution of this algorithm has been extensively studied before (Hong and Kung, 1981; Kwasniewski et al., 2019). Hong and Kung established a lower bound of $\Omega\left(\frac{m_{1}\cdot m_{2}\cdot m_{3}}{\sqrt{r}}\right)$ on $\texttt{OPT}_{RBP}$ using $S$-partitions. We show that the same lower bound also carries over to partial computations.
@@
The claim can be shown by combining the original lower bound proof of (Hong and Kung, 1981) with our $S$-edge partition concept. In the corresponding DAG, we have $m_{1}\cdot m_{2}$ sources, $m_{2}\cdot m_{3}$ sinks, and $m_{1}\cdot m_{2}\cdot m_{3}$ internal nodes. Each internal node has exactly one outgoing edge; let us call these internal edges. We consider where the $m_{1}\cdot m_{2}\cdot m_{3}$ internal edges can be sorted in an $S$-edge partition. The edge-terminal set of each $E_{i}$ contains at most $S$ sink nodes, and the edge-dominator set of $E_{i}$ contains at most $S$ source nodes. The proof in (Hong and Kung, 1981) shows that in the matrix multiplication DAG, for any subsets of sources $V_{1}$ and sinks $V_{2}$ with $|V_{1}|\leq S$, $|V_{2}|\leq S$, there are at most $O(S^{3/2})$ internal nodes on the paths from nodes in $V_{1}$ to nodes in $V_{2}$. Thus any set of $S$ sources can be a dominator for at most $O(S^{3/2})$ internal nodes, and hence also an edge-dominator for at most $O(S^{3/2})$ internal edges in $E_{i}$. As for internal nodes in the edge-dominator of $E_{i}$, we can have at most $S$ of them, each covering at most one internal edge. As such, $E_{i}$ contains at most $O(S^{3/2})+S$ internal edges. Therefore, any $S$-edge partitioning has at least $\Omega\left(\frac{m_{1}\cdot m_{2}\cdot m_{3}}{S^{3/2}+S}\right)$ classes; this is a lower bound on $\texttt{MIN}_{edge}(S)$. Selecting $S=2r$ and applying Theorem 6.5, the claim follows. ∎
@@ 6.3.3. Flash Attention
Due to the importance of the transformer architecture in machine learning, the I/O-efficient implementation of its crucial component, the self-attention mechanism, has been exhaustively studied. Recent work (Saha and Ye, 2024) in this area also used RBP to devise an I/O lower bound that matches the prominent Flash Attention algorithm. While attention consists of several steps, its main bottleneck is essentially matrix multiplication step $Q\cdot K^{T}$, where $Q$ and $K$ are both matrices of size $m\cdot d$. Due to this, a lower bound of $\Omega\left(\frac{m^{2}\cdot d}{\sqrt{r}}\right)$ follows from matrix multiplication. The work of (Saha and Ye, 2024) extends this with a stronger lower bound of $\Omega\left(\frac{m^{2}\cdot d^{2}}{r}\right)$ in the so-called large cache regime when $r\geq d^{2}$, which is indeed often the relevant case in practice. We again adapt this bound to PRBP.
@@
The first term follows from matrix multiplication as in (Saha and Ye, 2024), and carries over to PRBP due to Theorem 6.10. The second term is only smaller in the large cache regime, so assume $r\geq d^{2}$. Similarly to matrix multiplication, we can adapt the proof of (Saha and Ye, 2024) by shifting the analysis to internal edges instead of internal nodes. On a high level, consider an $S$-edge partition with $S=2r$, and focus on the internal edges in $E_{i}$. We have at most $2r$ internal nodes in the edge-dominator of $E_{i}$, covering at most $2r$ internal edges in $E_{i}$. The remaining internal edges must be dominated by source nodes. Assume that these remaining edges go towards $c$ sink nodes altogether, and these $c$ entries in $Q\cdot K^{T}$ span $c_{i}$ rows in $Q$ and $c_{j}$ columns in $K^{T}$. Then the edge-dominator of $E_{i}$ must contain $(c_{i\!}+_{\!}c_{j})_{\!}\cdot_{\!}d$ source nodes. Note that $c\!\leq\!c_{i}\!\cdot\!c_{j}$ means $\sqrt{c}\leq c_{i}\!+\!c_{j}$, and hence $(c_{i}\!+\!c_{j})\!\cdot\!d\leq 2r$ implies $c\leq\frac{4\cdot r^{2}}{d^{2}}$. Since their in-degree is $d$, the $c$ sink nodes can be the endpoints of at most $\frac{4\cdot r^{2}}{d}$ internal edges in $E_{i}$. Together with the $2r$ internal edges mentioned earlier, and using $r\geq d^{2}$, we get that $E_{i}$ has at most $O\left(\frac{r^{2}}{d}\right)$ internal edges. With altogether $m^{2}\!\cdot_{\!}d$ internal edges in the DAG, we get that $\texttt{MIN}_{edge}(2r)\geq\Omega\left(\frac{m^{2}\cdot d^{2}}{r^{2}}\right)$, and the bound follows from Theorem 6.5. ∎
@@ 7. Finding the optimal pebbling
Finally, we discuss the complexity of finding the best pebbling strategy in PRBP. In particular, we show that it is NP-hard to even approximate $\texttt{OPT}_{PRBP}$ to any non-trivial factor. The same hardness result has been shown for the original RBP in recent work (Böhnlein et al., 2025); however, in order to extend this proof technique to PRBP, we need to modify the gadgets used in their construction.
@@
The construction of (Böhnlein et al., 2025) shows that it is already NP-hard to distinguish between an I/O cost of $2$ and an I/O cost of $n^{1-\varepsilon}$. The main ingredients in this construction are the so-called level gadgets; we need modify these so that they still ensure the same desired properties in PRBP when we have partial computations. With the level gadgets adjusted, the rest of the construction can remain unchanged, proving the same hardness result for PRBP. For more details on the original proof, we refer the reader to (Böhnlein et al., 2025).
@@
Figure 5. Adding an auxiliary level (dotted grey part) to the level gadgets of the construction in (Böhnlein et al., 2025). Auxiliary levels always have the same size as the original level above them.
@@ 8.1. Alternative model variants
Except for the one-shot property, we defined and analyzed RBP according to its original definition of Hong and Kung (Hong and Kung, 1981), which is the most well-known version of the game. However, different variants of RBP were also studied before. We briefly outline how these relate to PRBP, and discuss more details in Appendix B.
If we remove the one-shot restriction, we obtain a model where the same node may be computed multiple times, which can be used sometimes to save I/O costs. This variant has been studied by Hong and Kung and others (Hong and Kung, 1981; Papp and Wattenhofer, 2020). The main modeling challenge here is that re-computation could be defined in multiple different ways in the partial-computing setting. Note that allowing re-computation can have a different effect on our theorems: for instance, Theorem 4.8 carries over easily to this case, Proposition 4.7 can be adapted with minor changes, whereas adapting e.g. the tools in Section 6 is more challenging.
Another variant of RBP also allows a different compute step which ‘slides’ a red pebble to $v$ from one of its inputs. The sliding model aims to capture in-place computations; this is somewhat similar to PRBP, which also assumes that the inputs of $v$ are aggregated into $v$ in an in-place fashion. However, in general, partial computations are a notably more powerful tool for saving I/O costs; in particular, our Proposition 4.7 is easy to adapt to the case when $\texttt{OPT}_{RBP}$ is understood in the sliding model.
In another extension towards practice, we also assign a small constant cost to the compute steps (Papp and Wattenhofer, 2020; Blocki et al., 2018). In the one-shot model, this only adds a fixed cost to any pebbling; however, this does change the ratio between optimal costs, and hence affects e.g. Theorem 7.1. Note that in order to compare $\texttt{OPT}_{RBP}$ and $\texttt{OPT}_{PRBP}$ here, one would need to define how the node-based compute costs in RBP are translated into edge-based partial compute costs in PRBP.
@@ 8.2. Practical remarks and future directions
While the usage of partial computations to reduce I/O cost has not been formalized and studied in general DAGs earlier, the general idea has of course been applied before (directly or indirectly) in several applications. For instance, we have seen how partial computations can reduce the I/O cost in a matrix-vector multiplication. In the case of dense matrix-matrix multiplication, the same approach is known as the outer-product formulation and e.g. used as the inner microkernel to state-of-the-art methods such as BLIS (Willenbring, 2015). In a broader sense, our results also indicate that e.g. in generalized linear algebraic programs, the distributive property of semirings can be exploited for practical gains; the same fact has been evidenced before when exploiting re-computation to tile through successive operations (e.g. (Demmel et al., 2008; McColl and Tiskin, 1999)).
We find that one particularly promising direction for future work is the version of the PRBP model extended with re-computations. This model incorporates two very natural methods to reduce the I/O cost of a computation in practice, and hence it would be essential to understand its general properties, and to analyze the cost of linear algebraic or other computations in this model. Another natural step towards practice is to generalize the PRBP model to computations that have both associative and non-associative operations, e.g. allowing for partial computations as long as a given order of the inputs is respected.
@@ Appendix A Further proof details
The steps of a RBP strategy with $r=4$ that only uses $3$ I/O steps is as follows: load $u_{0}$; compute $u_{1}$; delete $u_{0}$; compute $w_{1}$; compute $w_{2}$; compute $w_{3}$; delete $w_{1}$; delete $w_{2}$; compute $w_{4}$; delete $w_{3}$; delete $u_{1}$; load $u_{0}$; compute $u_{2}$; delete $u_{0}$; compute $v_{1}$; compute $v_{2}$; delete $w_{4}$; delete $u_{2}$; compute $v_{0}$; save $v_{0}$.
@@ A.2. Binary and $k$-ary trees
In RBP, with the exception of the bottom two levels, again for every node $v$ we have $(k-1)$ in-neighbors whose values will need to be saved and reloaded later to free up pebbles and compute the last subtree of $v$. Altogether, the number of non-trivial I/O operations in the optimal pebbling becomes
@@
In PRBP, on the other hand, partial computations allow us to compute the bottom $r=k+1$ levels without any I/O steps. Every node $v$ above these levels will again require $2\cdot(k-1)$ I/O steps, regardless of whether we compute $v$ partially in several iterations, or if we save $(k-1)$ of its in-neighbors and load them all later. The number of non-trivial I/O steps here is altogether
@@
Given a graph $G_{0}$ on $n$ nodes, if all nodes have degree $(n-1)$, then the entire graph is a clique, and we are finished. Otherwise, we can invoke $\mathcal{A}$ for each node of $G_{0}$ separately. If there are any nodes $v$ in $G_{0}$ that are not included in any maximum clique, we can remove these nodes (and their incident edges). This leaves all maximum cliques intact, so we can then recursively apply our algorithm on the remaining graph. On the other hand, if all nodes are included in a maximum clique, then consider an arbitrary node $v$ with degree strictly less than $(n-1)$. This implies that there is another node $u$ in $G_{0}$ that is not adjacent to $v$. Note that $u$ is also part of a maximum clique, and this does not include $v$; as such, even after removing $v$, at least one of the maximum cliques in the graph remains intact. Hence we can again remove $v$ (and its incident edges), and invoke our algorithm on the remaining graph. ∎
As discussed, the NP-hardness of maxinset-vertex and maxclique-vertex are equivalent, so Lemma A.1 follows naturally.
@@
Below we discuss some further details for the proof of Theorem 4.8. Note that the main idea of the original construction is analyzed in more detail in (Papp and Wattenhofer, 2020; Böhnlein et al., 2025).
Firstly, we note that the original construction of (Papp and Wattenhofer, 2020) does not initially use the gadget of Proposition 4.6 to form the groups $H_{1}(u_{0})$ and $H_{2}(u_{0})$, but instead just takes the same source nodes, and draws an edge from all of them to a single node. In RBP this already ensures that we need to have a red pebble on all the nodes of the gadget simultaneously. The pebble collection gadget is then only used to adapt the construction to a setting where the indegree of the DAG is bounded by a small constant. In contrast to this, in PRBP, the use of the gadget is essential; with partial computations, adding just a singe target node would not guarantee that we need to have a red pebble on all the inputs simultaneously.
@@
The main idea of selecting the chain size $\ell$ large enough in the gadgets has already been outlined. In practice, the proof is a bit more technical with two different length parameters $\ell$ and $\ell_{0}$. The total length of the chain will be $\ell$, and the chain will consist of an initial part of length $2(r-2)$, then a long part of length $\ell_{0}$, then a part of length $n_{0}$, and then another long part of length $\ell_{0}$. We ensure that any of the long parts is already sufficient in itself to guarantee suboptimality if we do not place all pebbles in the gadget, i.e. ł$\frac{\l_{0}}{2(r-1)}$ is already higher than the cost of any reasonable solution that pebbles each gadget in one go. Altogether, $\ell=2\ell_{0}+n_{0}+2(r-2)$.
@@
As for the rest of our parameters, we can select $b$ as a large constant; it has to be larger than $|Z_{1}|=|Z_{2}|=3$ in order to ensure that the I/O gain from the consecutive visitation of a gadget-pair is still much larger than the worst-case cost of computing $w$, and hence regardless of $w$’s role, the optimal pebbling will still always follow a maximum independent set. This defines $r=b+4n_{0}+5$ as discussed above. Finally, if we pebble each gadget simultaneously, but cannot visit any pair consecutively, we get a non-trivial I/O cost of at most $n_{0}\cdot b$ for reloading the $b$ merged nodes in the gadget-pairs, $2\cdot|E_{0}|$ for having to reload the middle nodes in the chains (where $|E_{0}|$ is the number of edges in $G_{0}$), plus $6$ for reloading $Z_{1}$ and $Z_{2}$. We need to ensure that $\frac{\ell_{0}}{2(r-2)}$ is larger than the sum of this. Furthermore, we add another $r-1$ here, to ensure that even the non-trivial I/O cost incurred by the gadget is larger than this sum. Thus altogether, we want to guarantee
@@
which can be satisfied with a choice of $\ell_{0}=\Theta(r\cdot(n_{0}+|E_{0}|+r))=\Theta(n_{0}^{2}+n_{0}\cdot|E_{0}|)$. Recall that then $\ell=2\ell_{0}+n_{0}+2(r-2)$, which is in the same magnitude.
@@
We then discuss further details for the proof of Theorem 7.1. Recall that in order to adapt this proof, we only add a given number of auxiliary levels before each original level in the construction of (Böhnlein et al., 2025). The remaining part of the proof (the towers formed from the gadgets, and the edges connecting them) is unchanged.
Recall that a level gadget is a chain $(u_{1},...,u_{\ell})$ with $(u_{i},u_{i+1})\!\in\!E$ for $i\!\in\![\ell\!-\!1]$. If the level $(u_{1},...,u_{\ell})$ is followed by the level $(v_{1},...,v_{\ell^{\prime}})$, then $(u_{i},v_{i})\!\in\!E$ for $i\!\in\![\min(\ell,\ell^{\prime})]$. If $\ell\!>\!\ell^{\prime}$, then also $(u_{i},v_{\ell^{\prime}})\!\in\!E$ for $\ell^{\prime}\!<\!i\!\leq\!\ell$. For an illustration of the gadgets, see Figure 3 in the full version of (Böhnlein et al., 2025). Our auxiliary levels are simply extra levels between the original levels of the construction, which always have the size of the next original level that follows. Adding these new levels does not affect the behavior in RBP at all: we can still transition between the original levels in the same way as before, just with more intermediate steps. As long as the total number of auxiliary levels is polynomial in the DAG size $n$, whole reduction remains polynomial.
In contrast to RBP, if have have a level $(u_{1},...,u_{\ell})$ in PRBP, it is possible that the red pebble from e.g. the node $u_{1}$ is already removed before we place a red pebble on $u_{\ell}$, and as such, it is not easy to show that an optimal solution will always pebble a level in a consecutive fashion. Thus in the analysis for PRBP, we will instead say that our strategy is on a given original level $(u_{1},...,u_{\ell})$ of a tower if we already placed a red pebble on all of $u_{1},...,u_{\ell}$ at least once, but the same does not hold for the following level.

### v2

@@ 6.3.2. Matrix Multiplication
Matrix multiplication is a fundamental operation in computer science, and also an essential building block in various applications. Given two matrices of sizes $m_{1}\times m_{2}$ and $m_{2}\times m_{3}$, respectively, we consider the standard matrix multiplication algorithm which executes $m_{1}\!\cdot\!m_{2}\!\cdot\!m_{3}$ multiplications. The I/O-efficient execution of this algorithm has been extensively studied before (RBpebbling1; RBpebbling6). Hong and Kung established a lower bound of $\Omega\left(\frac{m_{1}\cdot m_{2}\cdot m_{3}}{\sqrt{r}}\right)$ on $\texttt{OPT}_{RBP}$ using $S$-partitions. We show that the same lower bound also carries over to partial computations.
@@
The claim can be shown by combining the original lower bound proof of (RBpebbling1) with our $S$-edge partition concept. In the corresponding DAG, we have $m_{1}\cdot m_{2}$ sources, $m_{2}\cdot m_{3}$ sinks, and $m_{1}\cdot m_{2}\cdot m_{3}$ internal nodes. Each internal node has exactly one outgoing edge; let us call these internal edges. We consider where the $m_{1}\cdot m_{2}\cdot m_{3}$ internal edges can be sorted in an $S$-edge partition. The edge-terminal set of each $E_{i}$ contains at most $S$ sink nodes, and the edge-dominator set of $E_{i}$ contains at most $S$ source nodes. The proof in (RBpebbling1) shows that in the matrix multiplication DAG, for any subsets of sources $V_{1}$ and sinks $V_{2}$ with $|V_{1}|\leq S$, $|V_{2}|\leq S$, there are at most $O(S^{3/2})$ internal nodes on the paths from nodes in $V_{1}$ to nodes in $V_{2}$. Thus any set of $S$ sources can be a dominator for at most $O(S^{3/2})$ internal nodes, and hence also an edge-dominator for at most $O(S^{3/2})$ internal edges in $E_{i}$. As for internal nodes in the edge-dominator of $E_{i}$, we can have at most $S$ of them, each covering at most one internal edge. As such, $E_{i}$ contains at most $O(S^{3/2})+S$ internal edges. Therefore, any $S$-edge partitioning has at least $\Omega\left(\frac{m_{1}\cdot m_{2}\cdot m_{3}}{S^{3/2}+S}\right)$ classes; this is a lower bound on $\texttt{MIN}_{edge}(S)$. Selecting $S=2r$ and applying Theorem 6.5, the claim follows. ∎
@@ 6.3.3. Flash Attention
Due to the importance of the transformer architecture in machine learning, the I/O-efficient implementation of its crucial component, the self-attention mechanism, has been exhaustively studied. Recent work in this area also used RBP to devise an I/O lower bound that matches the prominent Flash Attention algorithm (ml1).
While attention consists of several steps, its main bottleneck is essentially a matrix multiplication step $Q\cdot K^{T}$, where $Q$ and $K$ are both matrices of size $m\times d$. However, in contrast to Theorem 6.10, it is a vital difference here that the entries of $Q\cdot K^{T}$ are not sink nodes, so we may not need to save them to slow memory. Due to this, the I/O complexity of the problem is more intricate. When we have $r\leq d^{2}$, a lower bound of $\Omega\left(\frac{m^{2}\cdot d}{\sqrt{r}}\right)$ follows easily from matrix multiplication: intuitively, with $\frac{m^{2}\cdot d}{\sqrt{r}}\geq m^{2}$, saving the matrix $Q\cdot K^{T}$ of size $m_{\!}\times_{\!}m$ to slow memory has little impact on the total cost. However, in the so-called large cache regime when $r\geq d^{2}$, which is often the relevant case in practice, the Flash Attention algorithm (flashattention) can obtain a lower I/O complexity of $O\left(\frac{m^{2}\cdot d^{2}}{r}\right)$. The work of (ml1) shows a matching lower bound; we again extend this lower bound to PRBP.
@@
When $r\leq d^{2}$, the first term in the minimum is smaller, and the problem can be reduced to matrix multiplication as in (ml1). The bound then carries over to PRBP due to Theorem 6.10.
The second term is only smaller in the large cache regime, so assume $r\geq d^{2}$. In this case, similarly to Theorem 6.10, we can adapt the proof of (ml1) by shifting the analysis to internal edges from internal nodes. The computational DAG of attention is described in detail in (ml1). The relevant part of the DAG is the matrix multiplication, which is represented by $2_{\!}\cdot_{\!}m_{\!}\cdot_{\!}d$ source nodes (entries in $Q$ and $K^{T}$), $m^{2}_{\!}\cdot d$ internal nodes that have two of these sources as inputs, and $m\cdot m$ nodes that represent an entry in $Q\cdot K^{T}$ (we call these root nodes) and have incoming edges from $d$ internal nodes. Each root node also has an outgoing edge to the remaining part of the DAG (specifically, to a next node that exponentiates the value). We again refer to the edges from internal nodes to roots as internal edges, and we call a root node and its in-neighbors together an internal tree. Note that internal trees are disjoint.
On a high level, consider an $S$-edge partition of this DAG with $S=2r$, and focus on the internal edges in each class $E_{i}$. Both the edge-dominator and the edge-terminal set of $E_{i}$ has size at most $2r$, so there are at most $4r$ internal trees that contain a node form either of these sets. These trees contain at most $4r\!\cdot\!d$ internal edges altogether.
If an internal tree is not among these $4r$ trees, but it still has an internal edge that is contained in $E_{i}$, then we will call this an extra tree. Since an extra tree contains no nodes from the edge-terminal set of $E_{i}$, the outgoing edge from its root node must also be in $E_{i}$. The paths starting from a source and containing this edge can only be covered by the source itself, since the extra tree contains no nodes of the edge-dominator set. As such, all the $2d$ source nodes which have a path to this root must be in the edge-dominator of $E_{i}$.
Assume that we have $c$ extra trees, and that the $c$ entries in $Q\cdot K^{T}$ corresponding to their root nodes altogether span $c_{i}$ rows in $Q$ and $c_{j}$ columns in $K^{T}$. Then the edge-dominator of $E_{i}$ must contain all the $(c_{i\!}+_{\!}c_{j})_{\!}\cdot_{\!}d$ source nodes in these rows and columns. Note that $c\!\leq\!c_{i}\!\cdot\!c_{j}$ implies $\sqrt{c}\leq c_{i}\!+\!c_{j}$. Due to this, $(c_{i}\!+\!c_{j})\!\cdot\!d\leq 2r$ implies $c\leq\frac{4\cdot r^{2}}{d^{2}}$. The root nodes have in-degree $d$, so the $c$ extra trees contain at most $\frac{4\cdot r^{2}}{d}$ internal edges in $E_{i}$. Together with the $4r\cdot d$ internal edges mentioned earlier, and using $r\geq d^{2}$, we get that $E_{i}$ has at most $O\left(\frac{r^{2}}{d}\right)$ internal edges. With altogether $m^{2}\!\cdot_{\!}d$ internal edges in the DAG, we get that
$\texttt{MIN}_{edge}(2r)\geq\Omega\left(\frac{m^{2}\cdot d^{2}}{r^{2}}\right)\,$,
and then the bound follows from Theorem 6.5. ∎
@@ 7. Finding the optimal pebbling
Finally, we discuss the complexity of finding the best pebbling strategy in PRBP. In particular, we show that it is NP-hard to even approximate $\texttt{OPT}_{PRBP}$ to any non-trivial factor. The same hardness result has been shown for the original RBP in recent work (mpp); however, in order to extend this proof technique to PRBP, we need to modify the gadgets used in their construction.
@@
The construction of (mpp) shows that it is already NP-hard to distinguish between an I/O cost of $2$ and an I/O cost of $n^{1-\varepsilon}$. The main ingredients in this construction are the so-called level gadgets; we need modify these so that they still ensure the same desired properties in PRBP when we have partial computations. With the level gadgets adjusted, the rest of the construction can remain unchanged, proving the same hardness result for PRBP. For more details on the original proof, we refer the reader to (mpp).
@@
Figure 5. Adding an auxiliary level (dotted grey part) to the level gadgets of the construction in (mpp). Auxiliary levels always have the same size as the original level above them.
@@ 8.1. Alternative model variants
Except for the one-shot property, we defined and analyzed RBP according to its original definition by Hong and Kung (RBpebbling1), which is the most well-known version of the game. However, different variants of RBP were also studied before. We briefly outline how these relate to PRBP, and discuss this in more detail in Appendix B.
If we remove the one-shot restriction, we obtain a model variant where the same node may be computed multiple times. Such re-computation steps can sometimes be used to save I/O costs, and hence they were allowed in the work of Hong and Kung and some others (RBpebbling1; RBpebbling3). When adapting this variant to PRBP, the main modeling challenge is that re-computation could be defined in multiple different ways in a partial-computing setting. We also note that allowing re-computation can have a rather different effect on our results: for instance, Theorem 4.8 carries over easily to this case, Proposition 4.7 can be adapted with minor changes, whereas adapting e.g. the tools in Section 6 is more challenging.
In another variant of RBP, we also allow a different compute step which ‘slides’ a red pebble to $v$ from one of its inputs. The sliding model aims to capture in-place computations; this is somewhat similar to PRBP, which also assumes that the inputs of $v$ are aggregated into $v$ in an in-place fashion. However, in general, partial computations allow us to save significantly more I/O cost than these sliding steps. In particular, Proposition 4.7 on the large difference between $\texttt{OPT}_{RBP}$ and $\texttt{OPT}_{PRBP}$ is also easy to adapt to the case when $\texttt{OPT}_{RBP}$ is understood in the sliding model.
In another extension towards practice, we can also assign a small constant cost to the compute steps (RBpebbling3; RBpebbling8). In the one-shot model, this only adds a fixed cost to any pebbling; however, this already affects the ratio between optimal costs in e.g. Theorem 7.1. Note that in order to compare $\texttt{OPT}_{RBP}$ and $\texttt{OPT}_{PRBP}$ in this variant, we would need to define how the node-based compute costs in RBP are translated into edge-based partial compute costs in PRBP.
@@ 8.2. Practical remarks and future directions
While the usage of partial computations to reduce I/O cost has not been formalized and studied in general DAGs earlier, the general idea has of course been applied before (directly or indirectly) in several applications. For instance, we have seen how partial computations can reduce the I/O cost in a matrix-vector multiplication. In the case of dense matrix-matrix multiplication, the same approach is known as the outer-product formulation and e.g. used as the inner microkernel to state-of-the-art methods such as BLIS (aj3). In a broader sense, our results also indicate that e.g. in generalized linear algebraic programs, the distributive property of semirings can be exploited for practical gains; the same fact has been evidenced before when exploiting re-computation to tile through successive operations (e.g. (aj2; aj1)). However, on the negative side, the inapproximability result shows that the problem is very challenging for irregular sparse computations in general.
We find that one promising direction for future work is the version of the PRBP model extended with re-computations. This model incorporates two very natural methods to reduce the I/O cost of a computation in practice, and hence it would be essential to understand its general properties, and to analyze the cost of linear algebraic or other computations in this model.
Note that another natural generalization of PRBP is to model computations that have both associative and non-associative operations, e.g. allowing for partial computations as long as a given order of the inputs is respected. However, such a setting can also be captured easily with a modification of the input DAG (by splitting up some of the nodes), as discussed before in (partial).
@@ Appendix A Further proof details
The steps of an RBP strategy with $r=4$ that only uses $3$ I/O steps is as follows: load $u_{0}$; compute $u_{1}$; delete $u_{0}$; compute $w_{1}$; compute $w_{2}$; compute $w_{3}$; delete $w_{1}$; delete $w_{2}$; compute $w_{4}$; delete $w_{3}$; delete $u_{1}$; load $u_{0}$; compute $u_{2}$; delete $u_{0}$; compute $v_{1}$; compute $v_{2}$; delete $w_{4}$; delete $u_{2}$; compute $v_{0}$; save $v_{0}$.
@@ A.2. Binary and $k$-ary trees
In RBP, with the exception of the bottom two levels, again for every node $v$ we have (k1)(k-1) in-neighbors whose values will need to be saved and reloaded later to free up pebbles and compute the last subtree of $v$. Altogether, the number of non-trivial I/O operations in the optimal pebbling becomes
@@
In PRBP, on the other hand, partial computations allow us to compute the bottom $r=k+1$ levels without any I/O steps. Every node $v$ above these levels will again require $2\cdot(k-1)$ I/O steps, regardless of whether we compute $v$ partially in several iterations, or if we save (k1)(k-1) of its in-neighbors and load them all later. The number of non-trivial I/O steps here is altogether
@@
Given a graph $G_{0}$ on $n$ nodes, if all nodes have degree (n1)(n-1), then the entire graph is a clique, and we are finished. Otherwise, we can invoke $\mathcal{A}$ for each node of $G_{0}$ separately. If there are any nodes $v$ in $G_{0}$ that are not included in any maximum clique, we can remove these nodes (and their incident edges). This leaves all maximum cliques intact, so we can then recursively apply our algorithm on the remaining graph. On the other hand, if all nodes are included in a maximum clique, then consider an arbitrary node $v$ with degree strictly less than (n1)(n-1). This implies that there is another node $u$ in $G_{0}$ that is not adjacent to $v$. Note that $u$ is also part of a maximum clique, and this does not include $v$; as such, even after removing $v$, at least one of the maximum cliques in the graph remains intact. Hence we can again remove $v$ (and its incident edges), and invoke our algorithm on the remaining graph. ∎
As discussed, the NP-hardness of maxinset-vertex and that of maxclique-vertex are equivalent, so Lemma 4.10 follows naturally.
@@
Below we discuss some further details for the proof of Theorem 4.8. Note that the main idea of the original construction is analyzed in more detail in (RBpebbling3; mpp).
Firstly, we note that the original construction of (RBpebbling3) does not initially use the gadget of Proposition 4.6 to form the groups $H_{1}(u_{0})$ and $H_{2}(u_{0})$, but instead just takes the same source nodes, and draws an edge from all of them to a single node. In RBP this already ensures that we need to have a red pebble on all the nodes of the gadget simultaneously. The pebble collection gadget is then only used to adapt the construction to a setting where the in-degree of the DAG is bounded by a small constant. In contrast to this, in PRBP, the use of the gadget is essential; with partial computations, adding just a singe target node would not guarantee that we need to have a red pebble on all the inputs simultaneously.
@@
The main idea of selecting the chains large enough in the gadgets has already been outlined. The actual proof is a bit more technical with two length parameters $\ell$ and $\ell_{0}$. Each chain starts with an initial part of length (r2)(r-2) to ensure that all I/O steps in the remaining parts are non-trivial; then a long part of length $\ell_{0}$, then a part of length $n_{0}$, and then another long part of length $\ell_{0}$. We ensure that any of the long parts is already sufficient in itself to guarantee suboptimality if we do not place all pebbles in the gadget, i.e. $\frac{\ell_{0}}{2(r-1)}$ is already higher than the cost of any reasonable solution that pebbles each gadget in one go. Altogether, $\ell=2\ell_{0}+n_{0}+(r-2)$.
@@
As for the rest of our parameters, we can select $b$ as a large constant; it has to be larger than $|Z_{1}|=|Z_{2}|=3$ in order to ensure that the I/O gain from the consecutive visitation of a gadget-pair is still much larger than the worst-case cost of computing $w$, and hence regardless of $w$’s role, the optimal pebbling will still always follow a maximum independent set. This defines $r=b+4n_{0}+5$ as discussed above. Finally, if we pebble each gadget simultaneously, but cannot visit any pair consecutively, we get a non-trivial I/O cost of at most $n_{0}\cdot b$ for reloading the $b$ merged nodes in the gadget-pairs, $2\cdot|E_{0}|$ for having to reload the middle nodes in the chains (where $|E_{0}|$ is the number of edges in $G_{0}$), plus $6$ for reloading $Z_{1}$ and $Z_{2}$. We need to ensure that $\frac{\ell_{0}}{2(r-2)}$ is larger than the sum of this. Furthermore, we add another r1r-1 here, to ensure that even the non-trivial I/O cost incurred by the gadget is larger than this sum. Thus altogether, we want to guarantee
@@
which can be satisfied with a choice of $\ell_{0}=\Theta(r\cdot(n_{0}+|E_{0}|+r))=\Theta(n_{0}^{2}+n_{0}\cdot|E_{0}|)$. Recall that then $\ell=2\ell_{0}+n_{0}+(r-2)$, which is in the same magnitude.
@@
We then discuss further details for the proof of Theorem 7.1. Recall that in order to adapt this proof, we only add a given number of auxiliary levels before each original level in the construction of (mpp). The remaining part of the proof (the towers formed from the gadgets, and the edges connecting them) is unchanged.
Recall that a level gadget is a chain $(u_{1},...,u_{\ell})$ with $(u_{i},u_{i+1})\!\in\!E$ for $i\!\in\![\ell\!-\!1]$. If the level $(u_{1},...,u_{\ell})$ is followed by the level $(v_{1},...,v_{\ell^{\prime}})$, then $(u_{i},v_{i})\!\in\!E$ for $i\!\in\![\min(\ell,\ell^{\prime})]$. If $\ell\!>\!\ell^{\prime}$, then also $(u_{i},v_{\ell^{\prime}})\!\in\!E$ for $\ell^{\prime}\!<\!i\!\leq\!\ell$. For an illustration of the gadgets, see Figure 3 in the full version of (mpp). Our auxiliary levels are simply extra levels between the original levels of the construction, which always have the size of the next original level that follows. Adding these new levels does not affect the behavior in RBP at all: we can still transition between the original levels in the same way as before, just with more intermediate steps. As long as the total number of auxiliary levels is polynomial in the DAG size $n$, the whole reduction remains polynomial.
In contrast to RBP, if have have a level $(u_{1},...,u_{\ell})$ in PRBP, it is possible that the red pebble from e.g. the node $u_{1}$ is already removed before we place a red pebble on $u_{\ell}$, and as such, it is not easy to show that an optimal solution will always pebble a level in a consecutive fashion. Thus in the analysis for PRBP, we will instead say that our strategy is currently on a given original level $(u_{1},...,u_{\ell})$ of a tower if we already placed a red pebble on all of $u_{1},...,u_{\ell}$ at least once, but the same does not hold for the following level.